The weights of male leopards in a particular region are normally distributed with mean 55kg and standard deviation 6kg.
- (a)Find the probability that a randomly chosen male leopard from this region weighs between 46 and 62 kg.[4]
- (b)The weights of female leopards in this region are normally distributed with mean 42kg and standard deviation σkg. It is known that 25% of female leopards in the region weigh less than 36kg. Find the value of σ.[3]
- (c)The distributions of the weights of male and female leopards are independent of each other. A male leopard and a female leopard are each chosen at random. Find the probability that both the weights of these leopards are less than 46kg.[4]
Mark scheme
Marking points from the official mark scheme (Feb/Mar 2022). The mark or mark type is shown in brackets.
- M1 46 or 62, 55 and 6 substituted into ±standardisation formula once. Condone 6² and continuity correction ±0.5
- B1 Both standardisation values correct, accept unsimplified
- M1 Calculating the appropriate area from stated Φs of z-values, must be probabilities.
- A1 0.8115 < p < 0.812
- P(46 < X < 62) = P((46-55)/6 < Z < (62-55)/6)
- = P(-1.5 < Z < 7/6)
- = P(-1.5 < Z < 1.1666...)
- = Φ(1.1666...) - Φ(-1.5)
- = Φ(1.1666...) - (1 - Φ(1.5))
- = 0.8784 + 0.9332 - 1 = 0.8116
- B1 CAO, critical z-value
- M1 36 and 42 substituted in ±standardisation formula, no continuity correction, not σ², √σ, equated to a z-value
- A1 WWW. Only dependent on M.
- P(F < 36) = 0.25
- P(Z < (36-42)/σ) = 0.25
- P(Z < -6/σ) = 0.25
- From tables, z = -0.674
- -6/σ = -0.674
- σ = 8.899 ≈ 8.90
- M1 FT value from part (a) or Correct: 1-Φ((46-55)/6), condone continuity correction, σ², √σ, and probability found. Condone unsupported correct value stated.
- M1 46, 42 and their 4(b) σ (or correct σ) substituted in ±standardisation formula, condone continuity correction, σ², √σ, and probability found Condone 4/their 8.90
- M1 Product of their 2 probabilities (0 < both < 1) Not 0.25 or their final answer to 4(a) used.
- A1 0.0449 < p < 0.0450
- P(male < 46) = P(Z < (46-55)/6) = P(Z < -1.5) = 1 - Φ(1.5) = 1 - 0.9332 = 0.0668
- P(female < 46) = P(Z < (46-42)/σ) = P(Z < 4/8.90) = P(Z < 0.4495) = Φ(0.4495) = 0.6736
- P(both < 46) = 0.0668 × 0.6736 = 0.045009
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →